# Ding Chen1  Qingchen Yu211footnotemark: 1  Pengyuan Wang211footnotemark: 1  Wentao Zhang322footnotemark: 2  &Bo Tang2  Feiyu Xiong2  Xinchi Li1  Minchuan Yang1  Zhiyu Li [^2] footnotemark: 2

### v1

@@ Abstract
With the release of the o1 model by OpenAI, reasoning models adopting slow thinking strategies have gradually emerged. As the responses generated by such models often include complex reasoning, intermediate steps, and self-reflection, existing evaluation methods are often inadequate. They struggle to determine whether the LLM output is truly equivalent to the reference answer, and also have difficulty identifying and extracting the final answer from long, complex responses. To address this issue, we propose xVerify, an efficient answer verifier for reasoning model evaluations. xVerify demonstrates strong capability in equivalence judgment, enabling it to effectively determine whether the answers produced by reasoning models are equivalent to reference answers across various types of objective questions. To train and evaluate xVerify, we construct the VAR dataset by collecting question-answer pairs generated by multiple LLMs across various datasets, leveraging multiple reasoning models and challenging evaluation sets designed specifically for reasoning model assessment. A multi-round annotation process is employed to ensure label accuracy. Based on the VAR dataset, we train multiple xVerify models of different scales. In evaluation experiments conducted on both the test set and generalization set, all xVerify models achieve overall F1 scores and accuracy exceeding 95%. Notably, the smallest variant, xVerify-0.5B-I, outperforms all evaluation methods except GPT-4o, while xVerify-3B-Ib surpasses GPT-4o in overall performance. These results validate the effectiveness and generalizability of xVerify. All resources for xVerify are available at https://github.com/IAAR-Shanghai/xVerify.
@@ 1 Introduction
With the emergence of chain of thought (CoT) prompting [35], researchers began to explicitly encourage LLMs to generate intermediate reasoning steps, thereby enhancing their ability to handle complex tasks. Following this, OpenAI introduced the o1 model [15], which proposed the concepts of slow thinking and scaling at test time. Specifically, the model is trained to output a detailed reasoning process before generating a final answer, significantly improving its performance on complex tasks. Inspired by this paradigm, a variety of reasoning models have emerged, such as DeepSeek-R1 [3] trained with GRPO, OpenAI’s o3-mini [26], and QwQ-32B [31]. However, the rise of reasoning models poses substantial challenges for evaluation. Since the outputs of these models often contain lengthy reasoning processes—potentially including redundant information, intermediate results, and even self-contradictions—it becomes significantly more difficult for evaluation tools to extract the final answer from such responses [2].
Developing evaluation methods tailored for LLM responses involving complex reasoning has become a key research focus. LLM reasoning is typically categorized into commonsense, logical, multi-hop, and mathematical reasoning [8]. Existing evaluation methods fall into automatic and human evaluation [2], with automatic evaluation gaining prominence due to its scalability and lower cost. The main automatic approaches for evaluating reasoning models include rule-based evaluation frameworks [13, 5, 27, 9, 25] and LLM-based judgment methods [20, 7, 18]. However, both approaches face limitations in reasoning model evaluation. Rule-based frameworks often struggle to extract final answers from lengthy reasoning traces, rely on strict formatting (e.g., syntactically correct LaTeX), and typically ignore the reasoning process itself—an oversimplification challenged by many researchers [36, 33, 14, 32]. Judge models are usually not optimized for reasoning evaluation and mainly produce qualitative scores or comments [7], making them more suitable for subjective questions. Objective tasks, in contrast, require accurate binary classification. Currently, effective automatic methods specifically designed for evaluating reasoning on objective questions remain lacking.
@@
To address these challenges, we introduce xVerify, an efficient LLM-answer verifier tailored for evaluating LLM responses to objective questions. xVerify processes the full LLM output, enabling it to accurately identify final answers from complex reasoning traces. It also supports robust equivalence checking, including symbol conversion (e.g., ‘alpha‘ → ‘$\alpha$‘), mathematical expression matching, and semantic alignment in natural language. Moreover, it is tolerant of formatting errors such as malformed LaTeX, making it applicable to a wide range of tasks, including math problems, multiple-choice, short-answer, and classification questions. To train and evaluate xVerify, we construct the Verify Answer for Reasoning (VAR) dataset, which includes responses from 19 LLMs across 24 reasoning benchmarks. All labels are verified through multi-round GPT-4o and human review. The dataset covers advanced reasoning models and benchmarks like GPQA, LiveMathBench, and AIME 2024. We fine-tune xVerify on a variety of base models (e.g., Qwen2.5, LLaMA, Gemma 2) and scales (0.5B–32B). Remarkably, even the smallest variant (xVerify-0.5B-I) surpasses existing evaluation methods—including 32B-sized models—on all metrics, while larger variants achieve F1 and accuracy over 95% on both test and generalization sets.
The main contributions of this paper are summarized as follows:
@@
  • We propose xVerify, an efficient answer verifier for reasoning model evaluations, and release multiple fine-tuned versions of xVerify. The checkpoints are publicly available [^1] Hugging Face collections: https://huggingface.co/collections/IAAR-Shanghai/xverify.
  • We conduct comprehensive comparative evaluations against multiple existing evaluation frameworks and judge models on both test and generalization datasets, thoroughly validating the effectiveness and applicability of xVerify.
@@ 2 Related Work
Evaluation methods have always been a crucial component in the development of LLM [2]. However, the open-ended nature of LLM outputs makes it difficult to apply standardized metrics, limiting the effectiveness of traditional evaluation methods [20]. The rise of reasoning models [26, 3, 31], which often generate lengthy and complex reasoning, further complicates evaluation. For objective tasks, the main challenge is to accurately extract the final answer from the LLM’s semi-structured output and compare it with the reference answer. Existing approaches are typically divided into human evaluation and automatic evaluation. While human evaluation offers flexibility, automatic methods are more cost-efficient and consistent [2]. Current automatic methods mainly include rule-based evaluation frameworks and LLM-based judgment methods.
@@
Rule-based methods are widely used in automatic evaluation frameworks such as LM Eval Harness [5], OpenCompass [27], UltraEval [9], and OpenAI Evals [25]. Tools like Math-Verify [13] also follow this approach, extracting final answers using regular expressions (RegEx) and comparing them with reference answers. However, LLM outputs often contain final answers in varied surface forms—e.g., "alpha" vs. "$\alpha$", "A" vs. "a", or "1000" vs. "$10^{3}$"—which can be semantically equivalent but textually different. While some tools support limited transformations, they typically handle only LaTeX expressions or simple string patterns, and struggle with basic semantic equivalence like "one hundred" vs. "100". For reasoning models, the output is usually lengthy and involves complex reasoning steps with intermediate results. This makes it difficult for regular expressions to accurately identify the final answer, causing rule-based approaches to frequently fail in such contexts. Moreover, prior work has shown that LLMs may revise or overturn their initial predictions during extended reasoning processes, exhibiting a kind of self-reflection [32]. At the same time, rule-based methods typically ignore the reasoning process and only evaluate the final answer, which has drawn criticism from many researchers—especially in the context of reasoning models [36, 33, 14]. Thus, rule-based evaluations have limited applicability in reasoning scenarios.
@@
LLM-based judgment methods use fine-tuned LLMs to evaluate the quality of other LLMs’ responses. Compared to traditional evaluation methods, they offer greater task adaptability, generate interpretable results, reduce evaluation costs, and can be applied across the LLM lifecycle [20, 7, 18]. For objective questions, these judge models can extract final answers from responses with intermediate reasoning or self-reflection. In recent years, many LLM-based judge models have emerged, including JudgeLM [39], PandaLM [34], Auto-J [21], Prometheus 2 [17], CompassJudger [1], CritiqueLLM [16], and Themis [12]. Judge models typically support pointwise, pairwise, and listwise evaluations [20], and some also serve as reward models in reinforcement learning. However, most are designed to assign scores to LLM outputs, making them more suitable for subjective evaluations like helpfulness, reliability, or relevance. For objective questions that require binary decisions (“correct” or “incorrect”), these models are less effective. Although scores can be binarized using thresholds, this approach is unreliable, as the models are not explicitly trained for such tasks. Moreover, the current LLM-based critic models and PRMs (Process Reward Models) exhibit subpar performance when detecting errors in long chain-of-thought responses generated by reasoning models [10]. Thus, while judge model holds promise for evaluating reasoning models, they require targeted training.
@@ 3 Problem Definition
For the stage of extracting the final answer, given a response $r$ to question $q$, which may include intermediate reasoning and multiple candidate answers, we denote the extracted candidates as $\mathrm{A}(r)$. To identify the final answer, we define a scoring function $\mathrm{S}:\mathrm{A}(r)\times\mathrm{Q}\rightarrow\mathbb{R}$ that measures the relevance or suitability of each candidate $a\in\mathrm{A}(r)$ to $q$, and select the final answer using the extraction function:
$\varepsilon(q,r)=\arg\max_{a\in\mathrm{A}(r)}\mathrm{S}(a,q).$ (1)
@@
$\psi_{math}(a_{ref}^{math},a_{final}^{math})=\begin{cases}1&\text{if }\Phi_{\text{norm}}^{math}(a_{ref}^{math})=\Phi_{\text{norm}}^{math}(a_{final}^{math}),\\ 0&\text{otherwise}\end{cases}$ (2)
@@
$\psi_{sym}(a_{ref}^{sym},a_{final}^{sym})=\begin{cases}1&\text{if }\Phi_{\text{norm}}^{sym}(a_{ref}^{sym})=\Phi_{\text{norm}}^{sym}(a_{final}^{sym}),\\ 0&\text{otherwise}\end{cases}$ (4)
Based on the above components, we define a unified equivalence function $\psi$ to determine whether the final answer $a_{final}$ matches the reference answer $a_{ref}$ across different modalities. The definition is:
$\psi(a_{final},a_{ref})=\begin{cases}1,&\text{if }\psi_{math}(a_{final}^{math},a_{ref}^{math})=1\\ &\quad\land\ \psi_{nl}(a_{final}^{nl},a_{ref}^{nl})=1\\ &\quad\land\ \psi_{sym}(a_{final}^{sym},a_{ref}^{sym})=1;\\ 0,&\text{otherwise}\end{cases}$ (5)
@@
To summarize, the overall evaluation function $\mathrm{E}$ is defined as:
$\mathrm{E}(q,r,a_{ref})=\psi\big{(}\varepsilon(q,r),\ a_{ref}\big{)}$ (6)
where $q$ is the objective question, $r$ is the response generated by the LLM, and $a_{ref}$ is the corresponding reference answer.
@@ 4 Methodology
The xVerify training and evaluation pipeline includes three main stages: collecting LLM responses, VAR dataset construction, and xVerify judge pipeline (see Figure 1). We first gather question–response pairs from various LLMs across four types of objective questions, including complex, reasoning-intensive examples. To ensure accurate labels, we employ multiple rounds of annotation and rechecking using both GPT-4o and human annotators. We also apply data augmentation to increase the dataset’s diversity and complexity. Finally, we train xVerify models of different sizes on the VAR dataset to evaluate long, multi-step answers—cases that are often difficult for existing evaluation methods. Section 4.1 details the dataset construction, and Section 4.2 describes the training process.
@@ 4.1 VAR Dataset
xVerify is designed to assess the correctness of reasoning models’ responses on objective questions. However, current judge models are mostly trained on tasks such as scoring or reviewing, and reasoning models with lengthy responses have only recently emerged. As a result, there is currently no suitable dataset for training xVerify. To better train and evaluate xVerify, we constructed a dedicated dataset named Verify Answer for Reasoning (VAR). Examples from the VAR dataset are provided in Appendix B.3.
@@ 4.1.1 LLM Response Generation
To ensure the diversity and coverage of the dataset, we selected 19 mainstream LLMs and 24 frequently used multilingual datasets to generate and collect responses. To better simulate the answering patterns of reasoning models in common evaluation scenarios, the chosen LLMs include recently released models such as the DeepSeek-R1-Distill series [3] and QwQ-32B [31]. Most of the other LLMs also support context lengths exceeding 32k tokens, enabling them to produce answers with extended reasoning chains. The selected datasets include high-difficulty benchmarks commonly used for evaluating reasoning models, such as GPQA [28], AIME 2024 [24], MATH [11], and LiveCodeBench [23], which typically require multi-step reasoning and computation to solve. During data generation, we also retained some extremely long responses, such as those exceeding 6k characters in length. Detailed information on all LLMs and datasets is provided in Appendix A.
To train and evaluate xVerify more effectively, we grouped the 24 datasets into four types based on question and answer formats: multiple choice, math, short answer, and classification. Multiple choice questions offer several labeled options; math includes questions where answers are mathematical expressions (e.g., numbers, equations), including mathematics and physics problems; short answer questions expect brief natural language responses like names or dates, with no strict format constraints; classification tasks involve selecting the correct label, such as for sentiment or topic classification.
To reflect realistic evaluation settings and generate a diverse set of Q&A samples, we designed multiple prompt templates for guiding the LLMs in response generation. The prompt configurations vary along several dimensions: 0-shot vs. 5-shot, with or without CoT, and with or without answer format restrictions (restrict), resulting in eight distinct prompt types. Details of all prompt templates are provided in Appendix D.1.
In total, we generated 191,600 Q&A samples using the 19 LLMs and 24 evaluation sets, providing a rich and diverse sample pool for constructing the dataset.

### v2

# xVerify: Efficient Answer Verifier for Reasoning Model Evaluations

@@ Abstract
With the release of OpenAI’s o1 model, reasoning models that adopt slow-thinking strategies have become increasingly common. Their outputs often contain complex reasoning, intermediate steps, and self-reflection, making existing evaluation methods and reward models inadequate. In particular, they struggle to judge answer equivalence and to reliably extract final answers from long, complex responses. To address this challenge, we propose xVerify, an efficient answer verifier for evaluating reasoning models. xVerify shows strong equivalence judgment capabilities, enabling accurate comparison between model outputs and reference answers across diverse question types. To train and evaluate xVerify, we construct the VAR dataset, which consists of question–answer pairs generated by multiple LLMs across various datasets. The dataset incorporates multiple reasoning models and challenging evaluation sets specifically designed for reasoning assessment, with a multi-round annotation process to ensure label quality. Based on VAR, we train xVerify models at different scales. Experimental results on both test and generalization sets show that all xVerify variants achieve over 95% F1 score and accuracy. Notably, the smallest model, xVerify-0.5B-I, outperforms all evaluation methods except GPT-4o, while xVerify-3B-Ib surpasses GPT-4o in overall performance. In addition, reinforcement learning experiments using xVerify as the reward model yield an 18.4% improvement for Qwen2.5-7B compared with direct generation, exceeding the gains achieved with Math Verify as the reward. These results demonstrate the effectiveness and generalizability of xVerify. All xVerify resources are available on GitHub.
xVerify: Efficient Answer Verifier for Reasoning Model Evaluations
Ding Chen1thanks: Equal contribution. $\dagger$ Corresponding authors   Qingchen Yu211footnotemark: 1   Pengyuan Wang211footnotemark: 1   Mengting Hu3 Wentao Zhang422footnotemark: 2Zhengren Wang5Bo Tang2Feiyu Xiong2 Xinchi Li1Chao Wang6Mingchuan Yang1Zhiyu Li [^2] footnotemark: 2 1 China Telecom Research Institute  2 MemTensor (Shanghai) Technology Co., Ltd. 2 College of Software, Nankai University  4 Center for Data Science, Peking University 5 Peking University  6 Data Development Center of China Telecom wentao.zhang@pku.edu.cn, lizy@iaar.ac.cn
@@ 1 Introduction
With the emergence of chain of thought (CoT) prompting (cot_2022_nips_google), researchers began to explicitly encourage LLMs to generate intermediate reasoning steps, thereby enhancing their ability to handle complex tasks. Following this, OpenAI introduced the o1 model (o1_2024_arXiv_openai), which proposed the concepts of slow thinking and scaling at test time. Specifically, the model is trained to output a detailed reasoning process before generating a final answer, significantly improving its performance on complex tasks. Inspired by this paradigm, a variety of reasoning models have emerged, such as DeepSeek-R1 (deepseek_r1_25_deepseek) trained with GRPO, OpenAI’s o3-mini (o3_mini_25_openai), and QwQ-32B (qwq32b_25_qwen). However, the rise of reasoning models has posed substantial challenges for evaluation. Because the reasoning processes output by these models may include redundant information, intermediate results, and self-reflections, current evaluation methods, as well as reward models in reinforcement learning, often prove ineffective (survey_2024_arXiv_Microsoft).
Developing scalable and accurate evaluation methods for LLMs on complex reasoning tasks (e.g., commonsense, logical, multi-hop, and mathematical reasoning) has become increasingly important (eval_survey_2023_arXiv_tianjin). While human annotation remains the gold standard, it is labor-intensive and difficult to scale. Automatic methods fall into two main categories: rule-based frameworks (opencompass_24_github; ultraeval_24_Tsinghua; math_verify_24_github; evals_24_github_OPENAI), which extract answers through strict formatting and pattern matching, and LLM-based judge models (llm_judge_survey_2024_arXiv_Tsinghua; llm_judge_survey_2025_arXiv_IDEA; llm_judge_survey_2025_arXiv_Arizona), which provide qualitative assessments or scores (survey_2024_arXiv_Microsoft). Both types of methods are commonly used to evaluate the performance of final models or to serve as reward models in reinforcement learning. However, rule-based methods struggle with diverse output formats and lengthy chains of thought; for example, Math-Verify (math_verify_24_github), adopted in the Open-R1111https://github.com/huggingface/open-r1 project, can only handle mathematical results that are strictly formatted and in fixed positions. Although judge models offer adaptability, they are not explicitly trained for objective ’correct/incorrect’ decision-making (llm_judge_survey_2025_arXiv_IDEA). Consequently, a robust, automated solution specifically tailored for objective reasoning evaluation is still lacking.
@@
To address these challenges, we introduce xVerify, an efficient LLM-answer verifier tailored for evaluating LLM responses to objective questions. xVerify processes the full LLM output, enabling it to accurately identify final answers from complex reasoning traces. It also supports robust equivalence checking, including symbol conversion (e.g., alpha → $\alpha$), mathematical expression matching, and semantic alignment in natural language. Moreover, it is tolerant of formatting errors such as malformed LaTeX, making it applicable to a wide range of tasks, including math problems, multiple-choice, short-answer, and classification questions. To train and evaluate xVerify, we construct the Verify Answer for Reasoning (VAR) dataset, which includes responses from 19 LLMs across 24 reasoning benchmarks. All labels are verified through multi-round GPT-4o and human review. The dataset covers advanced reasoning models and benchmarks like GPQA, LiveMathBench, and AIME 2024. We fine-tune xVerify on a variety of base models (e.g., Qwen2.5, LLaMA, Gemma 2) and scales (0.5B–32B). Remarkably, even the smallest variant (xVerify-0.5B-I) surpasses existing evaluation methods—including 32B-sized models—on all metrics, while larger variants achieve F1 and accuracy over 95% on both test and generalization sets. Furthermore, we conduct reinforcement learning (RL) experiments with xVerify as the reward model. Compared with direct generation, it shows an improvement of 18.4% for Qwen2.5-7B. This also represented a greater improvement than when Math Verify is used as the reward model. For Llama3.1-8B, we achieve similar improvements.
The main contributions of this paper can be summarized in three key points:
@@
  • We propose xVerify, an efficient answer verifier for evaluating reasoning models, and have released several fine-tuned versions that are publicly available on Hugging Face.
  • We comprehensively evaluate xVerify in two key capacities: as a judge model, demonstrating superior accuracy and robustness against existing methods on both in-domain and out-of-distribution benchmarks; and as a reward model in RL, where it improves policy performance over direct generation.
@@ 2 Related Work
Evaluation methods are a crucial component in the development of LLM (survey_2024_arXiv_Microsoft). However, the open-ended nature of LLM outputs makes it difficult to apply standardized metrics, limiting the effectiveness of traditional evaluation methods (llm_judge_survey_2024_arXiv_Tsinghua). The rise of reasoning models (o3_mini_25_openai; deepseek_r1_25_deepseek; qwq32b_25_qwen), generating lengthy and complex reasoning, further complicates evaluation. For objective tasks, the main challenge is to accurately extract the final answer from the LLM’s semi-structured output and compare it with the reference answer. Existing approaches are typically divided into human evaluation and automatic evaluation. While human evaluation offers flexibility, automatic methods are more cost-efficient and consistent (survey_2024_arXiv_Microsoft). Current automatic methods mainly include rule-based evaluation frameworks and LLM-based judgment methods.
@@
Rule-based methods are widely used in automatic evaluation frameworks such as LM Eval Harness (lm_eval_harness_21_Zenodo), OpenCompass (opencompass_24_github), UltraEval (ultraeval_24_Tsinghua), and OpenAI Evals (evals_24_github_OPENAI). Tools like Math-Verify (math_verify_24_github) also follow this approach, extracting final answers using regular expressions (RegEx) and comparing them with reference answers. However, LLM outputs often contain final answers in varied surface forms—e.g., "alpha" vs. $\alpha$", "A" vs. "a", or "1000" vs. $10^{3}$"—which can be semantically equivalent but textually different. While some tools support limited transformations, they typically handle only LaTeX expressions or simple string patterns, and struggle with basic semantic equivalence like "one hundred" vs. "100". For reasoning models, the output is usually lengthy and involves complex reasoning steps with intermediate results. This makes it difficult for regular expressions to accurately identify the final answer, causing rule-based approaches to frequently fail in such contexts. Moreover, prior work has shown that LLMs may revise or overturn their initial predictions during extended reasoning processes, exhibiting a kind of self-reflection (FTEvsTOE_24_arxiv_LMU). Additionally, rule-based methods typically ignore the reasoning process and only evaluate the final answer, which has drawn criticism from many researchers—especially in the context of reasoning models (cotper_2022_nips_google; self_consistency_2023_iclr_google; chain_2024_acl_google). Thus, rule-based evaluations are limited in reasoning scenarios.
@@
LLM-based judgment methods use fine-tuned LLMs to evaluate the quality of other LLMs’ responses. Compared to traditional evaluation methods, they offer greater task adaptability, generate interpretable results, reduce evaluation costs, and can be applied across the LLM lifecycle (llm_judge_survey_2024_arXiv_Tsinghua; llm_judge_survey_2025_arXiv_IDEA; llm_judge_survey_2025_arXiv_Arizona). For objective questions, these judge models can extract final answers from responses with intermediate reasoning or self-reflection. In recent years, many LLM-based judge models have emerged, including JudgeLM (judgelm_25_icrl), PandaLM (pandalm_24_icrl_Peking), Auto-J (auto_j_24_iclr_JiaoTong), Prometheus 2 (prometheus_2_24_acl_KAIST), CompassJudger (CompassJudger_1_24_arXiv_Shanghai_AI), CritiqueLLM (critiquellm_24_acl_Tsinghua), and Themis (themis_24_emnlp_Peking). Judge models typically support pointwise, pairwise, and listwise evaluations (llm_judge_survey_2024_arXiv_Tsinghua), and some also serve as reward models in reinforcement learning. However, most are designed to assign scores to LLM outputs, making them more suitable for subjective evaluations like helpfulness, reliability, or relevance. For objective questions that require binary decisions (“correct” or “incorrect”), these models are less effective. Although scores can be binarized using thresholds, this approach is unreliable, as the models are not explicitly trained for such tasks. Moreover, the current LLM-based critic models and PRMs (Process Reward Models) exhibit subpar performance when detecting errors in long chain-of-thought responses generated by reasoning models (deltabench_2025_arXiv_alibaba). Thus, while judge model holds promise for evaluating reasoning models, they require targeted training.
@@ 3 Problem Definition
For the stage of extracting the final answer, given a response $r$ to question $q$, which may include intermediate reasoning and multiple candidate answers, we denote the extracted candidates as $\mathrm{A}(r)$. To identify the final answer, we define a scoring function $\mathrm{S}:\mathrm{A}(r)\times\mathrm{Q}\rightarrow\mathbb{R}$ that measures the relevance or suitability of each candidate $a\in\mathrm{A}(r)$ to $q$, and select the final answer using the extraction function: $\varepsilon(q,r)=\arg\max_{a\in\mathrm{A}(r)}\mathrm{S}(a,q).$
@@
[…]
@@
[…]
Based on the above components, we define a unified equivalence function $\psi$ to determine whether the final answer $a_{final}$ matches the reference answer $a_{ref}$ across different modalities. Defined as:
$\psi(a_{final},a_{ref})=\begin{cases}1,&\text{if }\psi_{math}(a_{final}^{math},a_{ref}^{math})=1\\ &\quad\land\ \psi_{nl}(a_{final}^{nl},a_{ref}^{nl})=1\\ &\quad\land\ \psi_{sym}(a_{final}^{sym},a_{ref}^{sym})=1,\\ 0,&\text{otherwise}\end{cases}$ (4)
@@
To summarize, the overall evaluation function $\mathrm{E}$ is defined as: $\mathrm{E}(q,r,a_{ref})=\psi\big(\varepsilon(q,r),\ a_{ref}\big)$, where $q$ is the objective question, $r$ is the response generated by the LLM, and $a_{ref}$ is the corresponding reference answer.
@@ 4 Methodology
The xVerify training and evaluation pipeline includes three main stages: collecting LLM responses, VAR dataset construction, and the xVerify judge pipeline (Figure 1). We first gather question–response pairs from various LLMs across four types of objective questions, including complex, reasoning-intensive examples. To ensure accurate labels, we employ multiple rounds of annotation and rechecking using both GPT-4o and human annotators. We also apply data augmentation to increase the dataset’s diversity and complexity. Finally, we train xVerify models of different sizes on the VAR dataset to evaluate long, multi-step answers—cases challenging for existing evaluation methods. Section 4.1 details the dataset construction, and Section 4.2 describes the training process.
@@ 4.1 VAR Dataset
xVerify is designed to assess the correctness of reasoning models’ responses on objective questions. However, current judge models are mostly trained on tasks such as scoring or reviewing, and reasoning models with lengthy responses have only recently emerged. Thus, no suitable dataset exists for training xVerify. To better train and evaluate xVerify, we constructed a dedicated dataset named Verify Answer for Reasoning (VAR). Examples from the VAR dataset are provided in Appendix B.3.
@@ 4.1.1 LLM Response Generation
To ensure the diversity and coverage of the dataset, we selected 19 mainstream LLMs and 24 frequently used multilingual datasets to generate and collect responses. To better simulate the answering patterns of reasoning models in common evaluation scenarios, the chosen LLMs include recent models such as the DeepSeek-R1-Distill series (deepseek_r1_25_deepseek) and QwQ-32B (qwq32b_25_qwen). Most of the other LLMs also support context lengths exceeding 32k tokens, enabling them to produce answers with extended reasoning chains. The selected datasets include high-difficulty benchmarks commonly used for evaluating reasoning models, such as GPQA (gpqa_2024_colm_New_York), AIME 2024 (AIME_2024), MATH (math_2021_nips_UC), and LiveCodeBench (livemathbench_2024_arXiv_shanghailab), which typically require multi-step reasoning and computation to solve. During data generation, we also retained some extremely long responses, such as those exceeding 6k characters in length. Details on all LLMs and datasets are in Appendix A.
To train and evaluate xVerify more effectively, we grouped the 24 datasets into four types based on question and answer formats: multiple choice, math, short answer, and classification. Multiple choice questions offer several labeled options; math includes questions where answers are mathematical expressions (e.g., numbers, equations) in mathematics and physics; short answer questions expect brief natural language responses like names or dates, with no strict format constraints; classification tasks involve selecting the correct label, such as for sentiment or topic classification.
To reflect realistic evaluation settings and generate a diverse set of Q&A samples, we designed multiple prompt templates for guiding the LLMs in response generation. The prompt configurations vary along several dimensions: 0-shot vs. 5-shot, with or without CoT, and with or without answer format restrictions (restrict), resulting in eight distinct prompt types. Details of all prompt templates are provided in Appendix E.1.
In total, we generated 191,600 Q&A samples using the 19 LLMs and 24 evaluation sets, providing a diverse sample pool for constructing the dataset.